OpenStudy (anonymous):

A spherical balloon is expanding. Given that the radius is increasing at the rate of 2 inches per minute, at what rate is the volume increasing when the radius is 5 inches?

OpenStudy (campbell_st):

ok... so you need \[\frac{dV}{dt} = \frac{dV}{dr} \times \frac{dr}{dt}\] if \[V = \frac{4}{3} \pi r^3\] what is \[\frac{dV}{dr} = ?\]

OpenStudy (campbell_st):

and you are given \[\frac{dr}{dt} = 2\] so \[\frac{dV}{dt} = \frac{dV}{dr} \times 2\] just substitute r =5 for find the rate at which the volume is changing with respect to time

OpenStudy (anonymous):

SO IS DV/DR=4pir^2?

OpenStudy (campbell_st):

thats correct so \[\frac{dV}{dt} = 4\pi r^2 \times 2\] now substitute r = 5 and the units for the rate are cubic inches per minute

OpenStudy (anonymous):

where do i substitute the 5....

OpenStudy (campbell_st):

replace r with 5

OpenStudy (campbell_st):

so you are looking at \[\frac{dV}{dt} = 4 \pi (5)^2 \times 2\]

OpenStudy (anonymous):

4pi50

OpenStudy (anonymous):

do i multiply 4*50?

OpenStudy (campbell_st):

well \[\frac{dV}{dt} = 4 \times \pi \times 25 \times 2 = 200\pi\]

OpenStudy (anonymous):

okay! so is the answer left as 200pi?

OpenStudy (anonymous):

cubic inches per minute?

OpenStudy (campbell_st):

yep... thats my best guess

OpenStudy (anonymous):

thank you!